Suppose the market demand and supply are described by the following information: Q = 200 - 2P + 4i and Q = - 50 + 3P (5)

Determine the equilibrium price and quantity of the above market when i=50.

b)

Plot a graph showing the above results. 


1
Expert's answer
2021-08-24T10:08:35-0400

"Q_D = 200 - 2P +(4*50)"

"Q_D = 400 -2P"

"400 -2P = 3P-50 | - 400 -3P"

"-5P = -450 | Div -5"

"E_P = 90"

"E_Q = 3*90-50=220"

The market price is 90 and quantity is 220